# CityWalker: Learning Embodied Urban Navigation from Web-Scale Videos

### v2

@@ 1 Introduction
  • We demonstrate training on web-scale data significantly improves navigation performance in real-world experiments, allowing embodied agents to handle the complexities of urban scenarios. To ensure reproducibility, we will open-source our code and data.
@@ 2 Related Work
Navigation in Simulation. The introduction of high-performance photorealistic simulators [4, 5, 6, 7] has significantly boosted navigation research in indoor household environments. Based on the form of goal specification, navigation tasks are usually categorized into point-goal [26, 27, 4, 8], image-goal [28, 29, 30, 31], object-goal [32, 33, 34], and vision-language navigation [35, 36, 37]. While tasks like image-goal and object-goal navigation focus on efficient goal-finding or instruction understanding, point-goal navigation has a more straightforward goal specification and emphasizes safe and efficient trajectory planning to the goal. Our problem formulation resembles point-goal navigation in that accurate waypoints (represented as GPS coordinates) are readily available from navigation tools. Recently, point-goal navigation has been considered a ”solved” problem due to near-perfect performance achieved in simulators [8]. However, our investigation shows that waypoint-goal navigation in public urban spaces remains an unsolved challenge.
@@
In-the-Wild Video Learning. Learning from web-scale data has proven successful in language and vision tasks [19, 20, 46, 21]. While some works utilize in-the-wild videos to extract intermediate representations, priors, or reward functions [47, 48], a key challenge remains: the lack of action labels. Patel et al. [49] use a reconstruction pipeline to obtain hand-object trajectories as action labels for learning robot manipulation. Recent works also rely on prompting closed VLMs to get action labels [23]. LeLaN [24], a concurrent work most relevant to ours, also relies on VLM prompting and uses pretrained navigation models to generate action labels. In contrast, we process in-the-wild video data differently by using only the noisy pseudolabels from off-the-shelf visual SLAM tools [25]. This approach enables scalable and cost-effective action label generation for imitation learning in urban navigation tasks.
@@ 3.2 Evaluation Metrics
Figure 3: Evluation Metrics. Left: The orientation error is defined to be the angle between each predicted and ground truth action pair, labeled by $\theta_{n}$ in the figure. The action angle $\varphi_{\text{action}}$ and target angle $\varphi_{\text{target}}$ are defined with respect to the positive y-axis. Right: Both green and red trajectories are predicted actions. The green trajectory is the preferred one, having a large L2 distance but a small AOE. Vice versa for the red trajectory.
@@ 4 Experiment
Table 1: Benchmark on Offline Data. We evaluate three metrics in each critical scenario for all methods. Percentages under scenarios indicate their data proportions. The “Mean” column shows scenario means averaged over six scenarios; “All” shows sample means over all data samples. We highlight the best performance in bold and the second-best performance with underline.
@@
GNM [14] (zero-shot) $\downarrow$ AOE$(5)$ () 21.93 30.63 20.99 22.21 16.49 19.18 18.86 19.33
$\downarrow$ MAOE () 27.07 36.49 27.79 25.76 23.55 24.97 23.84 24.52
$\uparrow$ Arrival (%) 66.22 68.81 68.11 64.21 67.67 67.45 67.08 67.54
ViNT [44] (zero-shot) $\downarrow$ AOE$(5)$ () 11.45 27.49 8.31 9.84 9.58 6.79 6.70 7.64
$\downarrow$ MAOE () 17.52 31.82 16.10 14.21 16.80 13.63 12.59 13.63
$\uparrow$ Arrival (%) 69.07 67.83 69.36 70.08 66.97 69.61 70.54 70.15
ViNT [44] (fine-tune) $\downarrow$ AOE$(5)$ () 9.87 26.66 6.99 7.83 6.77 5.97 4.99 5.97
$\downarrow$ MAOE () 16.53 31.15 15.45 12.86 14.83 13.30 11.59 12.64
$\uparrow$ Arrival (%) 70.53 71.20 68.72 70.74 72.96 68.55 70.99 70.69
NoMaD [45] (zero-shot) $\downarrow$ AOE$(5)$ () 11.34 26.93 8.69 9.51 8.94 7.20 6.76 7.67
$\downarrow$ MAOE () 17.49 31.58 16.38 14.05 16.31 13.97 12.64 13.67
$\uparrow$ Arrival (%) 70.38 66.89 70.09 71.31 71.27 71.02 71.71 71.55
Ours (zero-shot) $\downarrow$ AOE$(5)$ () 8.43 17.88 6.19 10.68 6.64 5.22 3.98 5.09
$\downarrow$ MAOE () 16.54 26.76 15.46 16.29 16.25 13.10 11.36 12.66
$\uparrow$ Arrival (%) 79.09 69.31 71.44 78.82 84.05 84.51 86.42 84.12
Ours (fine-tune) $\downarrow$ AOE$(5)$ () 7.97 19.67 5.43 8.71 5.54 4.77 3.67 4.63
$\downarrow$ MAOE () 15.23 26.63 14.07 13.94 14.32 12.01 10.40 11.53
$\uparrow$ Arrival (%) 81.85 68.91 75.30 78.54 90.64 87.50 90.19 87.84
@@ 4.1 Setup
Data. We collected expert data via teleoperation for fine-tuning and offline testing. We collect data via a Unitree Go1 quadruped equipped with a Livox Mid-360 LiDAR and a webcam for RGB observations. We employed the LiDAR-SLAM method [57] to obtain the robot’s poses as ground-truth actions. Additionally, we retrieved location data from a smartphone using a webpage. We reduce noise from the quadruped’s vibrations by holding the phone [38]. In total, we gathered 15 hours of teleoperation data across various areas in New York City, allocating 6 hours for fine-tuning and 9 hours for testing.
@@ 4.2 Performance Benchmarking
Overall Benchmark. Table 1 shows the benchmark in different critical scenarios. Overall, our fine-tuned model has superior performance on all metrics in all scenarios except “detour”. We think the less competent performance in the detour scenario is due to the small portion of such data in our video training data. It can be verified by the fact that the fine-tuning performance gain is most significant in this scenario. It is worth noting that our zero-shot model has similar or even better performance compared to the fine-tuned ViNT model. This demonstrates the power of training with web-scale data, especially considering the domain and modality gap from human walking videos.
@@ 5 Conclusion
Limitation and Future Work. During real-world deployment, our method is sensitive to large GPS errors. This sensitivity results from our simple location data processing, which does not effectively handle significant positional inaccuracies. Future work will focus on enhancing the robustness of the model to location noise.
@@ B More Quantitative Results
VLM Performance. In Tab. III, we present the performance of the VLM (GPT-4o [61]) on our urban navigation tasks. Our results indicate that GPT-4o struggles to generate reasonable navigation actions off-the-shelf via prompting (see Fig. III). However, it performs reasonably in predicting the arrival status, likely because this sub-task is inherently more straightforward given the input of past and target locations.
Figure I: Data distribution. Top: The distribution of different weather and lightning condition in our video training data. Bottom: The distribution of different critical scenarios in our collected data.
@@ C More Qualitative Results
In Fig. IV, we provide more qualitative resting results on the offline data. We divide the results into three categories. Success: predicted action aligns well with ground truth action. Large error: predicted action does not align with ground truth but may still lead to successful navigation. Fail: predicted action may lead to failed navigation. The most significant observation is that large errors in offline data do not necessarily lead to failure in navigation, due to the multi-modality characteristic of policy learning. For example, in the fifth row, although the ground truth action takes a detour to the right of the traffic drum, the predicted action that goes straight from the left of the drum should also lead to successful navigation.
Figure III: VLM Prompt
Figure IV: More Qualitative Results. We provide more qualitative results in our offline testing. The results are categorized into success, large error, and fail. Success means the predicted action aligns with ground truth action. Large error mean prediction action does not align with ground truth but still lead to success navigation. Fails cases are those may lead to failed navigation.

### v3

@@ 1 Introduction
  • We demonstrate training on web-scale data significantly improves navigation performance in real-world experiments, allowing embodied agents to handle the complexities of urban scenarios. We also show model performance scales well with training data size.
@@ 2 Related Work
Navigation in Simulation. The introduction of high-performance photorealistic simulators [4, 5, 6, 7] has significantly boosted navigation research in indoor household environments. Based on the form of goal specification, navigation tasks are usually categorized into point-goal [26, 27, 4, 8], image-goal [28, 29, 30, 31], object-goal [32, 33, 34], and vision-language navigation [35, 36, 37]. While tasks like image-goal and object-goal navigation focus on efficient goal-finding or instruction understanding, point-goal navigation has a more straightforward goal specification and emphasizes safe and efficient trajectory planning to the goal. Our problem formulation resembles point-goal navigation in that accurate waypoints (represented as GPS coordinates) are readily available from navigation tools. Recently, point-goal navigation has been considered a "solved" problem due to near-perfect performance achieved in simulators [8]. However, our investigation shows that waypoint-goal navigation in public urban spaces remains an unsolved challenge.
@@
In-the-Wild Video Learning. Learning from web-scale data has proven successful in language and vision tasks [19, 20, 46, 21]. While some works utilize in-the-wild videos to extract intermediate representations, priors, or reward functions [47, 48], a key challenge remains: the lack of action labels. Patel et al. [49] use a reconstruction pipeline to obtain hand-object trajectories as action labels for learning robot manipulation. SelfD [50] argues BEV model is better than VO in pseudo-labeling, while we show VO works well with large-scale data. Recent works also rely on prompting closed VLMs to get action labels [23, 51]. LeLaN [24], a concurrent work most relevant to ours, also relies on VLM prompting and uses pretrained navigation models to generate action labels. In contrast, we process in-the-wild video data differently by using only the noisy pseudolabels from off-the-shelf visual SLAM tools [25]. This approach enables scalable and cost-effective action label generation for imitation learning in urban navigation tasks.
@@ 3.2 Evaluation Metrics
Figure 3: Evaluation Metrics. Left: The orientation error is defined to be the angle between each predicted and ground truth action pair, labeled by $\theta_{n}$ in the figure. The action angle $\varphi_{\text{action}}$ and target angle $\varphi_{\text{target}}$ are defined with respect to the positive y-axis. Right: Both green and red trajectories are predicted actions. The green trajectory is the preferred one, having a large L2 distance but a small AOE. Vice versa for the red trajectory.
@@
As orientation error might be noisy when action distance is small, we also use L2-distance as a complimentary evaluation metric.
@@ 4 Experiment
Table 1: Benchmark on Offline Data. We evaluate three metrics in each critical scenario for all methods. Percentages under scenarios indicate their data proportions. The “Mean" column shows scenario means averaged over six scenarios; “All" shows sample means over all data samples. We highlight the best performance in bold and the second-best performance with underline.
@@
GNM [14] (fine-tuned) $\downarrow$ L2 (m) 1.22 2.36 1.36 1.42 0.88 0.76 0.55 0.74
$\downarrow$ MAOE () 16.2 31.1 14.8 12.5 14.7 12.8 11.0 12.1
$\uparrow$ Arrival (%) 68.6 66.4 69.7 66.4 69.0 69.2 70.7 70.0
ViNT [44] (fine-tuned) $\downarrow$ L2 (m) 1.30 1.91 1.13 1.14 0.77 0.66 0.57 0.70
$\downarrow$ MAOE () 16.5 31.1 15.4 12.9 14.8 13.3 11.6 12.6
$\uparrow$ Arrival (%) 70.5 71.2 68.7 70.7 73.0 68.6 71.0 70.7
NoMaD [45] (fine-tuned) $\downarrow$ L2 (m) 1.39 2.49 1.56 1.55 1.06 0.95 0.76 0.74
$\downarrow$ MAOE () 19.1 35.1 18.5 15.6 18.1 14.3 12.8 12.1
$\uparrow$ Arrival (%) 68.6 66.4 69.7 66.4 69.0 69.3 70.7 70.0
Ours (zero-shot) $\downarrow$ L2 (m) 1.34 1.30 1.09 1.33 1.44 1.48 1.39 1.38
$\downarrow$ MAOE () 16.5 26.8 15.5 16.3 16.3 13.1 11.3 12.7
$\uparrow$ Arrival (%) 79.1 69.3 71.4 78.8 84.1 84.5 86.4 84.1
Ours (fine-tuned) $\downarrow$ L2 (m) 1.11 1.27 1.00 1.15 1.06 1.12 1.06 1.07
$\downarrow$ MAOE () 15.2 26.6 14.1 13.9 14.3 12.0 10.4 11.5
$\uparrow$ Arrival (%) 81.8 68.9 75.3 78.5 90.6 87.5 90.2 87.8
@@ 4.1 Setup
Data. We collected expert data via teleoperation for fine-tuning and offline testing. We collect data via a Unitree Go1 quadruped equipped with a Livox Mid-360 LiDAR and a webcam for RGB observations. We employed the LiDAR-SLAM method [59] to obtain the robot’s poses as ground truth actions. Additionally, we retrieved location data from a smartphone using a webpage. We reduce noise from the quadruped’s vibrations by holding the phone [38]. In total, we gathered 15 hours of teleoperation data across various areas in New York City, allocating 6 hours for fine-tuning and 9 hours for testing.
@@ 4.2 Performance Benchmarking
Overall Benchmark. Table 1 shows the benchmark in different critical scenarios. Overall, our fine-tuned model has superior performance on all metrics in all scenarios except “detour”. We think the less competent performance in the detour scenario is due to the small portion of such data in our video training data. It can be verified by the fact that the fine-tuning performance gain is most significant in this scenario. It is worth noting that our zero-shot model has similar or even better performance compared to the fine-tuned baselines. This demonstrates the power of large-scale pretraining, especially considering the domain and modality gap from human walking videos.
@@ 5 Conclusion
Limitation and Future Work. During real-world deployment, our current system implementation is sensitive to potential large location noises read naively from iPhone location services, which could be improved with better GPS hardware. Future work could also enhance the robustness of the model to such location noises.
@@ B More Quantitative Results
VLM Performance. In Tab. III, we present the performance of the VLM (GPT-4o [63]) on our urban navigation tasks. Our results indicate that GPT-4o struggles to generate reasonable navigation actions off-the-shelf via prompting. However, it performs reasonably in predicting the arrival status, likely because this sub-task is inherently more straightforward given the input of past and target locations.
Figure I: Data distribution. Top: The distribution of different weather and lightning conditions in our video training data. Bottom: The distribution of different critical scenarios in our collected data.
@@
Image Backbones. In Tab. IV, we show that our model performance is not sensitive to the choice of image backbones. This makes embodied depolyment very efficient. While our model with DiNOv2 backbone only has 1.7 fps inference speed on a RTX 3060 laptop, this can be boosted to 20 fps by switching to EfficientNetB0 backbone without sacrificing model performance.
@@ C More Qualitative Results
In Fig. III, we provide more qualitative resting results on the offline data. We divide the results into three categories. Success: predicted action aligns well with ground truth action. Large error: predicted action does not align with ground truth but may still lead to successful navigation. Fail: predicted action may lead to failed navigation. The most significant observation is that large errors in offline data do not necessarily lead to failure in navigation, due to the multi-modality characteristic of policy learning. For example, in the fifth row, although the ground truth action takes a detour to the right of the traffic drum, the predicted action that goes straight from the left of the drum should also lead to successful navigation.
Table IV: Comparison of backbones and architecture. All models are pretrained with 2000 hours of video and fine-tuned with expert data. Both metrics are taking the category mean. *Pretrained from ACO [65].
Metric EfficientNetB0 ResNet50 DiNOv2 ResNet34* ViNT**
MAOE ($\downarrow$) 15.33 15.16 15.23 15.13 15.26
L2 ($\downarrow$) 1.11 m 1.15 m 1.12 m 1.09 m 1.08 m
Figure III: More Qualitative Results. We provide more qualitative results in our offline testing. The results are categorized into success, large error, and fail. Success means the predicted action aligns with ground truth action. Large error mean prediction action does not align with ground truth but still lead to success navigation. Fails cases are those may lead to failed navigation.